y=-4/5+3 must have 3 linear points
I think one is (0,3) and another is (5,7)

Answer by rfer(16322)

did you leave the x out of equation?
y=-4/5x+3
(0,3) is one
--------------
0=-4/5x+3
0=-.8x+3
-.8x=-3
x=3.75
(3.75,0)
---------------
y=-.8(5)+3
y=-4+3
y=-1
(5,-1)
---------------
7=-.8x+3
4=-.8x
x=-5
(-5,7)
you lost track of the signs in part of your calculations